The following data were obtained in an experiment to determine the enthalpy of fusion of water (i.e. the amount of energy required to convert solid water to liquid water). Calculate the enthalpy of fusion of ice (in J/g) from the data. Assume the specific heat of liquid water is 4.184 J/g°C.

mass of water in calorimeter = 108.31 g
initial temperature of water = 25.1°C
mass of ice at 0°C = 5.45 g
final temperature of the water with melted ice = 19.5°C
1
Expert's answer
2020-05-12T11:34:45-0400

Heat lost by water = heat gained by ice

"m_1c\Delta T=m_2L+m_2c\Delta T"

"25.1\times4.184\times(25.1-19.5)=5.45\times L+5.45\times4.184\times19.5\\588.1=5.45\times L+444.65\\L=26.32J/g"